User Tools

Site Tools


handover:a1404

A1404

Hb:

  • Recording continued on HOB+0128/32000 from ~15369 GB (?) new number: 16.486 TB

Ke:

  • Recording continued on HOB+0123/32000 from ~14345 GB (?) new number: 1.7 TB
  • System check OK. Halt issued at 04:02:50 UT, schedule a1404 resumed at 04:23:13

Yg:

  • Recording continued on HOB+1006/16000 from ~10245 GB (?) new number: 5.6 TB
  • Wind stow 17:53 – 22:02.
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/a1404.txt · Last modified: 2014/09/07 22:13 by David Horsley